OpenStudy (anonymous):

find the sum of 21 + 12 + 6 + 3 + ... to 12 terms

OpenStudy (ybarrap):

Looks like the difference between each term is less than the previous by a factor of 3: 12=21-3*3, 6=12-3*2, etc...

OpenStudy (anonymous):

21+12+6+3+3+0-6-15-27-42-60-81 = -186 check my calculation for double proof

OpenStudy (anonymous):

@ybarrap How would you put that into the equation? We're using Sn = n/2 (2(u1)+(n-1)d

OpenStudy (ybarrap):

Let's work it out: 21 + (21 - 3*3) + (21 - 3*3 - 3*2) + (21 - 3*3 - 3*2 - 3*1) + ... This does not have a common difference and so is not an arithmetic series where you can use that formula. I've worked out a double sum of this and think that @helpme1.2 's method is easier and appears correct.